# Characterizations of voting rules based on majority margins

### v2

@@ 2.3 Proof of characterization
Next, for any $ab$ such that there are both designated $ab$-voters and also designated $ba$-voters in $V(\mathbf{P}^{+})$, by the second condition of Lemma 2.15 there must be designated $ab$$\top$-, $ab$$\bot$-, $ba$$\top$-, and $ba$$\bot$-voters in $V(\mathbf{P}^{+})$. Then we remove for each of these four categories a voter whose pre-image under $v$ has the largest third coordinate. Note that removing these four voters amounts to deleting two reversal pairs, since $G^{\top}_{ab}$ and $G^{\bot}_{ba}$ form a reversal pair and $G^{\bot}_{ab}$ and $G^{\top}_{ba}$ form a reversal pair. Thus, by Neutral Reversal it does not change the output of $F$. We continue doing so until there are either no designated $ab$-voters or no designated $ba$-voters. The second condition in Lemma 2.15 is preserved by the above operation of removing two pairs of voters, so we are guaranteed to hit one of these two cases.
@@
Call the resulting profile $\mathbf{P}^{\dagger}$. It is now clear that $\mathcal{M}_{\mathbf{P}^{\dagger}}$ is still $m$ and $\mathbf{P}^{\dagger}$ satisfies all the four conditions of Lemma 2.15. Thus, $\mathbf{P}^{\dagger}=\mathbf{D}_{v,m}$. Also, as we have only used Neutral Reversal and Preferential Compensation, we have $F(\mathbf{P}^{\dagger})=F(\mathbf{P}^{+})=F(\mathbf{P})$.
@@
Case 2: the parity of the number of voters in $\mathbf{P}$ and $\mathbf{Q}$ is odd. By Neutral Reversal, we may assume without loss of generality that some voter $i$ appears in both $\mathbf{P}$ and $\mathbf{Q}$ with the same ballot, so $\mathbf{P}_{i}=\mathbf{Q}_{i}$. Now for each pair of ballots $P,P^{\prime}$ in $\mathbf{P}$ that are not $\mathbf{P}_{i}$, proceed exactly as in the previous case. The resulting profile $\mathbf{P}^{\dagger}$ is identical to $\mathbf{P}_{i}+\mathbf{D}_{v,\mathcal{M}_{\mathbf{P}}-\mathcal{M}_{\mathbf{P}_{i}}}$ and hence
@@
Thus, we have replaced $P,P^{\prime}$ with a McGarvey pair. In other cases, more steps are needed to transform an initial pair into a reversal pair, which can then be eliminated. The process of adding McGarvey pairs may also create some quadruples of the form $G^{\top}_{ab},G^{\bot}_{ab},G^{\top}_{ba},G^{\bot}_{ba}$, which can be eliminated at the end. Ultimately we obtain the following profile, $\mathbf{D}_{v,\mathcal{M}_{\mathbf{P}}}$, where the number $k$ above a ranking $L$ indicates the number of voters—whose identities are determined by the function $v$—submitting the ranking $L$:
$2$ $2$ $1$ $1$ $2$ $2$ $1$ $1$ $2$ $2$ $1$ $1$
@@
Remark 2.17.
@@
A natural question is whether we can modify Theorem 2.9 by weakening Neutral Reversal to the axiom of Block Invariance (Holliday and Pacuit Forthcoming), which states that adding to a profile $\mathbf{P}$ exactly one copy of each linear order of $X(\mathbf{P})$ results in a profile $\mathbf{P}^{\prime}$ such that $F(\mathbf{P})=F(\mathbf{P}^{\prime})$. Unlike Neutral Reversal, Block Invariance is satisfied by the Plurality voting rule and IRV rule. However, the answer is that we cannot weaken Theorem 2.9 in this way. Fix some sequence of margin-based voting rules $F_{1},F_{2},\dots$. Then define a new voting rule $F$ as follows: if $|V(\mathbf{P})|$ is congruent to $k$ modulo $|X(\mathbf{P})|!$, for $k\leq|X(\mathbf{P})|!$, then $F(\mathbf{P})=F_{k}(\mathbf{P})$. Since each $F_{k}$ is margin-based, $F$ clearly satisfies Preferential Equality, whose definition does not involve a change in the number of voters. Moreover, adding to $\mathbf{P}$ a block of all linear orders of $X(\mathbf{P})$ results in a profile that is in the same congruence class modulo $|X(\mathbf{P})|!$ as $\mathbf{P}$ is and has the same margins, so $F$ satisfies Block Invariance. But it is easy to choose the sequence $F_{1},F_{2},\dots$ so that $F$ is not margin-based, because whether the number of voters in a profile $\mathbf{P}$ falls into a particular congruence class modulo $|X(\mathbf{P})|!$ is not uniquely determined by $\mathcal{M}_{\mathbf{P}}$.
@@ 3 Characterization for all profiles, modulo homogeneity
In this section, we consider the domain of all profiles, as well as a domain intermediate between linear profiles and all profiles. Say that a strict weak order on $X$ is a linear order with bottom indifference if whenever candidates $x$ and $y$ are ranked in a tie (i.e., neither $(x,y)\in\mathbf{P}(i)$ nor $(y,x)\in\mathbf{P}(i)$), then $x$ and $y$ are not ranked strictly above any candidate. Note that every linear order is considered a linear order with bottom indifference, even though it contains no indifference. Note also that the strict weak order where all candidates are tied together (represented by the empty set) is a linear order with bottom indifference as well.
@@
The importance of this domain is that in real political elections with ranked ballots, voters are not required to rank all the candidates. To turn such an election into one of our profiles, it is standard to consider all candidates unranked by a voter as being in a big tie below all candidates ranked by the voter. Thus, real political elections yield profiles in the $\mathsf{LOBI}$ domain.
@@ 3.1 Tiebreaking Compensation
Now that we are allowing ties in rankings (even if only at the bottom in LOBI), we introduce an axiom motivated by the idea that each voter should have equal tiebreaking power.
@@
A voting rule $F$ satisfies Tiebreaking Compensation if for all profiles $\mathbf{Q},\mathbf{Q}^{\prime}\in\mathrm{dom}(F)$, if in $\mathbf{Q}$ there are two voters $i$ and $j$ who each rank a set $Y$ of candidates in a tie, and $\mathbf{Q}^{\prime}$ is obtained from $\mathbf{Q}$ by $i$ replacing the tie $Y$ by some linear order $L$ of $Y$ and $j$ replacing the tie $Y$ by $L^{-1}$, then $F(\mathbf{Q})=F(\mathbf{Q}^{\prime})$.
@@
Let $\mathbf{P},\mathbf{P}^{\prime}$ be as in the definition of Preferential Compensation in Lemma 2.3, so $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by one voter switching $ab$ to $ba$ while another switches $ba$ to $ab$. Let $\mathbf{Q}$ be the profile exactly like $\mathbf{P}$ except that both voters rank $\{a,b\}$ in a tie (and the only tie since $\mathbf{P}$ has only linear ballots). Then setting $\mathbf{Q}^{\prime}:=\mathbf{P}$, we obtain $F(\mathbf{Q})=F(\mathbf{P})$ by Tiebreaking Compensation with $Y=\{a,b\}$. Similarly, by setting $\mathbf{Q}^{\prime}:=\mathbf{P}^{\prime}$, we obtain $F(\mathbf{Q})=F(\mathbf{P}^{\prime})$ by Tiebreaking Compensation. Hence $F(\mathbf{P})=F(\mathbf{P}^{\prime})$, as desired. Then $F$ also satisfies Preferential Equality by Lemma 2.3. ∎
@@
Let $F$ be the voting rule defined on LOBI as follows: if the profile contains a tie or two voters who rank different candidates in last place, then $F$ outputs all candidates; otherwise $F$ outputs the Plurality winners. Then $F$ satisfies Tiebreaking Compensation [^8] In the definition of Tiebreaking Compensation, $\mathbf{Q}$ has at least a tie, and since we assume here that $\mathbf{Q}\in\mathsf{LOBI}$, $\mathbf{Q}^{\prime}$ contains two voters who rank different candidates in last place, so $F(\mathbf{Q})=X(\mathbf{Q})=X(\mathbf{Q}^{\prime})=F(\mathbf{Q}^{\prime})$. but not Preferential Compensation or Preferential Equality (since Plurality does not satisfy these axioms).
@@
    The Pareto rule on the domain of all profiles satisfies Tiebreaking Compensation (and Preferential Compensation) but not Neutral Reversal.
@@
3.2 Homogeneity
In our characterizations for the domain of all profiles and for LOBI, we will add a weak assumption that is not satisfied by all margin-based voting rules but is normatively highly compelling. Fix a linear order of $\mathcal{V}$, which we call the alphabetic order of $\mathcal{V}$.
Definition 3.6.
Given a profile $\mathbf{P}$, let $2\mathbf{P}=\mathbf{P}+\mathbf{P}^{\prime}$ where $\mathbf{P}^{\prime}$ is a copy of $\mathbf{P}$ using the first $|V(\mathbf{P})|$-voters from $\mathcal{V}\setminus V(\mathbf{P})$ according to the alphabetic order of $\mathcal{V}$. A voting rule satisfies Homogeneity (and is called homogeneous) if for any profile $\mathbf{P}\in\mathrm{dom}(F)$, we have $F(\mathbf{P})=F(2\mathbf{P})$.
The original Homogeneity axiom of Smith 1973 states that $F(\mathbf{P})=F(n\mathbf{P})$ for any positive integer $n$, where $n\mathbf{P}$ has the obvious meaning, but we only need the doubling version in what follows. One of the only voting rules discussed in the literature that violates Homogeneity is the Dodgson rule (see Brandt 2009), but this is considered a pathology of the rule. It would be hard to justify to voters that if every voter brought a twin to the election who submitted the same ballot, this could change the election outcome.
Fact 3.7.
3.3 Proof of characterization
Theorem 3.8.
Let $F$ be a voting rule satisfying Homogeneity.
  1. 1.
    If the domain of $F$ is LOBI, then $F$ is margin-based if and only if $F$ satisfies Preferential Equality, Tiebreaking Compensation, and Neutral Reversal.
  2. 2.
    If the domain of $F$ is the domain of all profiles, then $F$ is margin-based if and only if $F$ satisfies Tiebreaking Compensation and Neutral Reversal.
Proof.
If a profile contains a ranking with a tie, duplicate the profile, so we have two copies of the ranking, and then linearize the previously tied part in opposite ways. By Homogeneity and Tiebreaking Compensation, this does not change the output of $F$. In this way, we can change any two initial profiles into linear profiles. Then we reason exactly as in the proof of Theorem 2.9 using Preferential Equality, which is available in part 2 thanks to Lemma 3.3.∎
Remark 3.9.
In Theorem 3.8, we can replace Neutral Reversal by the axiom of Neutral Indifference discussed in Section 4.2, according to which adding a voter with a fully indifferent ballot does not change the outcome. Clearly Neutral Indifference (applied twice) and Tiebreaking Compensation jointly imply Neutral Reversal.
In part 1 of Theorem 3.8, we can also replace Tiebreaking Compensation by what could be called Pure Tiebreaking Compensation: if two voters have the same rankings containing a tie $Y$, then changing the profile so that each voter linearizes $Y$ in opposite ways does not change the output of the voting rule. For in the proof of Theorem 3.8.1, we apply Tiebreaking Compensation to two copies of the same ranking. We could also replace Tiebreaking Compensation with Pure Tiebreaking Compensation in part 2 of Theorem 3.8, provided we add Preferential Equality, since unlike Tiebreaking Compensation, Pure Tiebreaking Compensation does not imply Preferential Equality for voting rules on the domain of all profiles. [^9] Given $Y\subsetneq X$ and rankings $P$ and $P^{\prime}$ of $X$, say that $P$ and $P^{\prime}$ agree outside $Y$, denoted $P\sim_{Y}P^{\prime}$, if (a) for all $x\in X\setminus Y$, either $xPy$ for all $y\in Y$ or $yPx$ for all $y\in Y$, and either $xP^{\prime}y$ for all $y\in Y$ or $yP^{\prime}x$ for all $y\in Y$, and (b) $P\setminus Y^{2}=P^{\prime}\setminus Y^{2}$. Let $\mathbb{P}$ be the set of profiles $\mathbf{P}$ such that there are two voters whose rankings $P$ and $P^{\prime}$ agree outside some $Y\subsetneq X(\mathbf{P})$ or are such that $P^{\prime}=P^{-1}$. Let $F$ be the voting rule such that if $\mathbf{P}\in\mathbb{P}$, then $F(\mathbf{P})=X(\mathbf{P})$, and otherwise $F(\mathbf{P})$ is the set of Plurality winners. Observe that if $\mathbf{Q}^{\prime}$ is obtained from $\mathbf{Q}$ by an application of Pure Tiebreaking Compensation, then $\mathbf{Q},\mathbf{Q}^{\prime}\in\mathbb{P}$, so $F$ satisfies Pure Tiebreaking Compensation. However, $F$ violates Preferential Equality for the profile $\mathbf{P}$ with just two ballots, $abcde$ and $dabec$, since switching $ab$ to $ba$ in the first ballot vs. the second ballot leads to profiles that are not in $\mathbb{P}$ and have different Plurality winners.
Fact 3.10.
4 Characterization modulo head-to-head
We now return to the distinction introduced in Section 1 between margin-based voting rules and the larger class of head-to-head voting rules. We already know as a consequence of Theorem 3.8.2 that a homogeneous, head-to-head voting rule on the domain of all profiles is margin-based if and only if it satisfies Tiebreaking Compensation and Neutral Reversal. However, once we make the assumption that the voting rule is head-to-head, we can provide an alternative axiomatization of margin-based rules, while also dropping the Homogeneity assumption. For this we need only two axioms, introduced in the following subsections.
In this section, we work with voting rules on the domain of all profiles (for LOBI, we refer the reader back to Theorem 3.8.1). For notational convenience, where $S_{1},\dots,S_{n}$ partition $X$,
$S_{1}>S_{2}>\dots>S_{n}$
4.1 Nonlinear Neutral Reversal
The first axiom we will use in our characterization of margin-based rules relative to the class of head-to-head rules is a strengthening of Neutral Reversal. It applies not only when two voters submit reversed linear orders but also when two voters submit reversed rankings with ties.
Definition 4.1.
Fact 4.2.
Let $F$ be the version of the Borda count for profiles of strict weak orders defined as follows: from a strict weak order $P$, a candidate $x$ receives |$|\{y\in X(\mathbf{P})\mid(x,y)\in P\}|$ points, and the candidates with the most points win. Then adding the pair of rankings $\{a\}>X(\mathbf{P})\setminus\{a\}$ and $X(\mathbf{P})\setminus\{a\}>\{a\}$ to a profile gives $|X(\mathbf{P})|-1$ points to $a$ and $1$ point to everyone else, which can change the set of winners, so $F$ does not satisfy Nonlinear Neutral Reversal. Yet it is easy to see that $F$ does satisfy Neutral Reversal.

### v3

@@ 2.3 Proof of characterization
Next, for any $xy$ such that there are both designated $xy$-voters and also designated $yx$-voters in $V(\mathbf{P}^{+})$, by the second condition of Lemma 2.15 there must be designated $\top-, xyxy-\bot$-, yxyx-$\top$-, and yxyx-$\bot$-voters in $V(\mathbf{P}^{+})$. Then we remove for each of these four categories a voter whose pre-image under $v$ has the largest third coordinate. Note that removing these four voters amounts to deleting two reversal pairs, since $G^{\top}_{xy}$ and $G^{\bot}_{yx}$ form a reversal pair and $G^{\bot}_{xy}$ and $G^{\top}_{yx}$ form a reversal pair. Thus, by Neutral Reversal it does not change the output of $F$. We continue doing so until there are either no designated $xy$-voters or no designated $yx$-voters. The second condition in Lemma 2.15 is preserved by the above operation of removing two pairs of voters, so we are guaranteed to hit one of these two cases. If all voters would be removed at the final step of this process, we first add the profile $\mathbf{D}_{v,0}$ fixed at the beginning before the final removal, which leaves the output of $F$ unchanged.
@@
Call the resulting profile $\mathbf{P}^{\dagger}$. It is now clear that $\mathcal{M}_{\mathbf{P}^{\dagger}}$ is still $m$. If the final addition of a reversal pair did not happen, then $\mathbf{P}^{\dagger}$ satisfies all the four conditions of Lemma 2.15, and thus, $\mathbf{P}^{\dagger}=\mathbf{D}_{v,m}$. If the final addition did happen, then $m$ must be the zero matrix since performing only margin-preserving operations led us to the empty “profile”, and by stipulation $\mathbf{P}^{\dagger}=\mathbf{D}_{v,m}$. Also, as we have only used Neutral Reversal and Preferential Compensation, we have $F(\mathbf{P}^{\dagger})=F(\mathbf{P}^{+})=F(\mathbf{P})$.
@@
Case 2: the parity of the number of voters in $\mathbf{P}$ and $\mathbf{Q}$ is odd. By Neutral Reversal, we may assume without loss of generality that some voter $i$ appears in both $\mathbf{P}$ and $\mathbf{Q}$ with the same ballot, so $\mathbf{P}_{i}=\mathbf{Q}_{i}$. Now for the part of the profile $\mathbf{P}$ that excludes the voter $i$, proceed exactly as in the previous case, including the final addition of the fixed reversal pair if $\mathcal{M}_{\mathbf{P}}-\mathcal{M}_{\mathbf{P}_{i}}$ is the zero matrix. Then after adding the voter $i$ back, the resulting profile $\mathbf{P}^{\dagger}$ is identical to $\mathbf{P}_{i}+\mathbf{D}_{v,\mathcal{M}_{\mathbf{P}}-\mathcal{M}_{\mathbf{P}_{i}}}$ and hence
@@
Below is an example showcasing how to reduce a profile $\mathbf{P}$ to its normal form $\mathbf{D}_{v,\mathcal{M}_{\mathbf{P}}}$ by adding and removing reversal pairs and applying preferential compensation.
@@
Thus, we have replaced $P,P^{\prime}$ with a McGarvey pair. In this case, the newly added voters end up having exactly the same ballots as the deleted $i_{3}$ and $i_{4}$. But this is necessary if we want to convert two profiles sharing the same margin matrix to exactly the same profile even using the same voters. In other cases, more steps are needed to transform an initial pair into a reversal pair, which can then be eliminated. For example, to turn $i_{5}$’s ballot into the reversal of $i_{6}$’s, we need to first swap $a$ with $c$ and then $d$ and then $b$, and then swap $d$ with $b$. These steps introduce $v(ca,\top,1)$ and $v(ca,\bot,1)$, $v(da,\top,1)$ and $v(da,\bot,1)$, $v(ba,\top,1)$ and $v(ba,\bot,1)$, and $v(bd,\top,2)$ and $v(bd,\bot,2)$ (since $v(bd,\top,1)$ and $(bd,\bot,1)$ have already been used). The process of adding McGarvey pairs may also create some quadruples of the form $G^{\top}_{xy},G^{\bot}_{xy},G^{\top}_{yx},G^{\bot}_{yx}$, which can be eliminated at the end. Ultimately we obtain the following profile, $\mathbf{D}_{v,\mathcal{M}_{\mathbf{P}}}$, where the number $k$ above a ranking $L$ indicates the number of voters—whose identities are determined by the function vv—submitting the ranking $L$:
22 22 11 11 22 22 11 11 22 22 11 11
@@
2.4 Weakening Neutral Reversal to Block Invariance
A natural question is whether we can modify Theorem 2.9 by weakening Neutral Reversal to the following axiom of Block Invariance (Holliday and Pacuit 2025; cf. Merlin 2003, Def. 4).
Definition 2.17.
A voting rule $F$ satisfies Block Invariance if for any $\mathbf{P},\mathbf{P}^{\prime}\in\mathrm{dom}(F)$ with $X(\mathbf{P})=X(\mathbf{P}^{\prime})$, if $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by adding, for each linear order $L$ of $X(\mathbf{P})$, exactly one voter submitting $L$ as their ranking, then $F(\mathbf{P})=F(\mathbf{P}^{\prime})$.
Unlike Neutral Reversal, Block Invariance is satisfied by the Plurality voting rule and IRV rule. However, the answer is that we cannot weaken Theorem 2.9 by replacing Neutral Reversal with Block Invariance, as shown by the following example.
Example 2.18.
@@ 2.4 Weakening Neutral Reversal to Block Invariance
Fix some sequence of margin-based voting rules $F_{0},F_{1},\dots$. Then define a new voting rule $F$ as follows: if $|V(\mathbf{P})|$ is congruent to $k$ modulo $|X(\mathbf{P})|!$, for $k\in\{0,\dots,|X(\mathbf{P})|!-1\}$, then $F(\mathbf{P})=F_{k}(\mathbf{P})$. Since each $F_{k}$ is margin-based, $F$ clearly satisfies Preferential Equality, whose definition does not involve a change in the number of voters. Moreover, adding to $\mathbf{P}$ a block of all linear orders of $X(\mathbf{P})$ results in a profile that is in the same congruence class modulo $|X(\mathbf{P})|!$ as $\mathbf{P}$ is and has the same margins, so $F$ satisfies Block Invariance. But it is easy to choose the sequence $F_{0},F_{1},\dots$ so that $F$ is not margin-based, because whether the number of voters in a profile $\mathbf{P}$ falls into a particular congruence class modulo $|X(\mathbf{P})|!$ is not uniquely determined by $\mathcal{M}_{\mathbf{P}}$.
Despite this example, if we avail ourselves of the normatively highly compelling axiom of Homogeneity (Smith 1973), we can indeed weaken Neutral Reversal to Block Invariance. For the definition of Homogeneity, fix a linear order of $\mathcal{V}$, which we call the alphabetic order of $\mathcal{V}$.
Definition 2.19.
Given a profile $\mathbf{P}$ and $n\geq 1$, let $n\mathbf{P}=\mathbf{P}+\mathbf{P}^{\prime}$ where $\mathbf{P}^{\prime}$ is n1n-1 copies of $\mathbf{P}$ using the first $(n-1)|V(\mathbf{P})|$ voters from $\mathcal{V}\setminus V(\mathbf{P})$ according to the alphabetic order of $\mathcal{V}$. A voting rule satisfies Homogeneity and is called homogeneous if for any profile $\mathbf{P}\in\mathrm{dom}(F)$, we have $F(\mathbf{P})=F(n\mathbf{P})$ for all $n\geq 2$.
One of the only voting rules discussed in the literature that violates Homogeneity is the Dodgson rule (see Brandt 2009), but this is considered a pathology of the rule. It would be hard to justify to voters that if every voter brought a twin to the election who submitted the same ballot, this could change the election outcome.
Fact 2.20.
Let us now prove that by assuming Homogeneity, we can weaken Neutral Reversal to Block Invariance in the characterization of margin-based rules. The proof is a simple modification of the proof using Neutral Reversal.
Theorem 2.21.
Let $F$ be a voting rule satisfying Homogeneity. Then $F$ is margin-based if and only if it satisfies Preferential Equality and Block Invariance.
Proof.
For the non-trivial direction, suppose we have profiles $\mathbf{P}$ and $\mathbf{Q}$ such that $\mathcal{M}_{\mathbf{P}}=\mathcal{M}_{\mathbf{Q}}$. Let $n=|X(\mathbf{P})|=|X(\mathbf{Q})|$, $\mathbf{P}^{\prime}=n!\mathbf{P}$, and $\mathbf{Q}^{\prime}=n!\mathbf{Q}$. Then again $\mathcal{M}_{\mathbf{P}^{\prime}}=\mathcal{M}_{\mathbf{Q}^{\prime}}$, but more importantly, both $|V(\mathbf{P}^{\prime})|$ and $|V(\mathbf{Q}^{\prime})|$ are multiples of $n!$. By Homogeneity, $F(\mathbf{P})=F(\mathbf{Q})$ iff $F(\mathbf{P}^{\prime})=F(\mathbf{Q}^{\prime})$, so it suffices to prove the latter to show that $F$ is margin-based.
Our first step is to make the voters in $\mathbf{P}^{\prime}$ and $\mathbf{Q}^{\prime}$ vote blocks of all possible linear orders of the candidates and then use Block Invariance to eliminate them. In our proof of Theorem 2.9, we used Neutral Reversal for this step. To use only Block Invariance, when introducing the two new designated $xy$-voters with the reversal rankings $yxL$ and $L^{-1}xy$, we also introduce new voters for all other linear orders of the candidates, so that a new block is introduced. Call this block a designated-$xy$-block. To formally track the identity of these added voters, we again use a $v$ function, now of the type $X^{(2)}\times\mathsf{Lin}(X)\times\mathbb{N}_{>0}\to\mathcal{V}\setminus(V(\mathbf{P}^{\prime})\cup V(\mathbf{Q}^{\prime}))$, where $\mathsf{Lin}(X)$ is the set of all linear orders on $X$. That is, we are replacing $\{\top,\bot\}$ with $\mathsf{Lin}(X)$. A designated-$xy$-block is then a set $\{v(xy,R,k)\mid R\in\mathsf{Lin}(X)\}$ for some $k$ (the $k$-th designated-xyxy-block). The voter $v(xy,R,k)$ will always vote $R$ for any $R\in\mathsf{Lin}(X)$ except when $R=yxL$, where again $L$ is the alphabetical order on $X\setminus\{x,y\}$. When introducing such a block without changing the outcome of the election, we let $v(xy,yxL,k)$ vote $yxL$. Then, after a use of Preferential Compensation (recall Lemma 2.3) to swap $xy$ to $yx$ in the ballot of an old voter, the voter $v(xy,yxL,k)$ votes $xyL$, and now these $n!$ voters in $\{v(xy,R,k)\mid R\in\mathsf{Lin}(X)\}$ form a profile whose margin matrix is equal to that of the profile with just the McGarvey pair $xyL$ and $L^{-1}xy$, i.e., with a margin of $2$ for $xy$, 2-2 for $yx$, and 0 for all other pairs. After all the old voters are eliminated, we then have to make sure that no designated-$xy$-blocks and designated-$yx$-blocks exist simultaneously. Observe that if we have both a designated-$xy$-block and a designated-$yx$-block, then they in fact form two blocks of all linear orders of the candidates, since $v(xy,yxL,k)$ and $v(yx,xyL,k^{\prime})$ now vote each other’s original ballot, namely $xyL$ and $yxL$, from before we used Preferential Compensation. Thus, they can be removed together without changing the outcome by Block Invariance. If all voters are removed, we then add back a fixed block of all linear orders not using any of the original voters in either $\mathbf{P}$ and $\mathbf{Q}$. In exactly the same way as we proved Lemma 2.15, we can prove that there is only one profile to realize a given non-zero margin matrix $m$ of even parity using these designated blocks in the order given by the third index $k$ without redundancy. Thus, $\mathbf{P}^{\prime}$ and $\mathbf{Q}^{\prime}$ can be turned into the same profile using only Preferential Compensation and Block Invariance, so $F(\mathbf{P}^{\prime})=F(\mathbf{Q}^{\prime})$, as desired. ∎
@@ 3 Characterization for all profiles, modulo homogeneity
In this section, we consider the domain of all profiles, as well as a domain intermediate between linear profiles and all profiles. Say that a strict weak order $P$ on $X$ is a linear order with bottom indifference if whenever candidates $x$ and $y$ are ranked in a tie (i.e., neither $(x,y)\in P$ nor $(y,x)\in P$), then $x$ and $y$ are not ranked strictly above any candidates. Note that every linear order is considered a linear order with bottom indifference, even though it contains no indifference. Note also that the strict weak order where all candidates are tied together (represented by the empty set) is a linear order with bottom indifference as well.
@@
The importance of this domain is that in real political elections with ranked ballots, voters are not required to rank all the candidates. To turn such an election into one of our profiles, it is standard to consider all candidates unranked by a voter as being in a big tie below all candidates ranked by the voter. Thus, real political elections yield profiles in the $\mathsf{LOBI}$ domain. Moreover, some voting rules used in real political elections, such as IRV, are traditionally defined only for the $\mathsf{LOBI}$ domain (though generalizations have been considered, e.g., in Delemazure and Peters 2024). Finally, as we will see (in Theorem 3.8), the axioms sufficient to force the use of a margin-based voting rule when the domain is assumed to be the domain of all profiles might not be sufficient to force the use of a margin-based voting rule when the domain is assumed to be the $\mathsf{LOBI}$ domain of real political elections, so for normative purposes, we must analyze the $\mathsf{LOBI}$ case separately.
@@ 3.1 Tiebreaking Compensation
Now that we are allowing ties in rankings (even if only at the bottom in $\mathsf{LOBI}$), we introduce an axiom motivated by the idea that each voter should have equal tiebreaking power.
@@
A voting rule $F$ satisfies Tiebreaking Compensation if for all profiles $\mathbf{Q},\mathbf{Q}^{\prime}\in\mathrm{dom}(F)$, if in $\mathbf{Q}$ there are two voters $i$ and $j$ and a set $Y$ of candidates which is an indifference class for both $i$ and $j$, and $\mathbf{Q}^{\prime}$ is obtained from $\mathbf{Q}$ by $i$ replacing the tie $Y$ by some linear order $L$ of $Y$ and $j$ replacing the tie $Y$ by $L^{-1}$, then $F(\mathbf{Q})=F(\mathbf{Q}^{\prime})$.
@@
Let $\mathbf{P},\mathbf{P}^{\prime}$ be as in the definition of Preferential Compensation in Lemma 2.3, so $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by one voter switching $xy$ to $yx$ while another switches $yx$ to $xy$. Let $\mathbf{Q}$ be the profile exactly like $\mathbf{P}$ except that both voters rank $\{x,y\}$ in a tie (and the only tie since $\mathbf{P}$ has only linear ballots). Then setting $\mathbf{Q}^{\prime}:=\mathbf{P}$, we obtain $F(\mathbf{Q})=F(\mathbf{P})$ by Tiebreaking Compensation with $Y=\{x,y\}$. Similarly, by setting $\mathbf{Q}^{\prime}:=\mathbf{P}^{\prime}$, we obtain $F(\mathbf{Q})=F(\mathbf{P}^{\prime})$ by Tiebreaking Compensation. Hence $F(\mathbf{P})=F(\mathbf{P}^{\prime})$, as desired. Then $F$ also satisfies Preferential Equality by Lemma 2.3. ∎
@@
Let $F$ be the voting rule defined on $\mathsf{LOBI}$ as follows: if the profile contains a tie or two voters who rank different candidates in last place, then $F$ outputs all candidates; otherwise $F$ outputs the Plurality winners. Then $F$ satisfies Tiebreaking Compensation [^11] In the definition of Tiebreaking Compensation, $\mathbf{Q}$ has at least a tie, and since we assume here that $\mathbf{Q}\in\mathsf{LOBI}$, $\mathbf{Q}^{\prime}$ contains two voters who rank different candidates in last place, so $F(\mathbf{Q})=X(\mathbf{Q})=X(\mathbf{Q}^{\prime})=F(\mathbf{Q}^{\prime})$. but not Preferential Compensation or Preferential Equality (since Plurality does not satisfy these axioms).
@@
    The strict Pareto rule on the domain of all profiles satisfies Tiebreaking Compensation (and Preferential Compensation) but not Neutral Reversal.